OpenStudy (anonymous):

(a+3)(a-2) use foil i tried but got the wrond answer please help me solve this

OpenStudy (anonymous):

multiply each thing in first bracket by each in the second one

OpenStudy (anonymous):

a*a+a*(-2)+3*a+3*(-2)= compute the rest yourself...

jimthompson5910 (jim_thompson5910):

First: a*a = a^2 Outer: a*(-2) = -2a Inner: 3*a = 3a Last: 3*(-2) = -6 Add up the terms a^2+(-2a)+3a+(-6) a^2+a-6 So the answer is a^2+a-6

OpenStudy (anonymous):

(a+3)(a-2) a^2-2a+3a-6 a^2+a-6

OpenStudy (anonymous):

steps... (a*a)+(a*-2)+(3*a)+(3*-6) understand?